OpenStudy (anonymous):

Please help me! The ratio of two numbers is 1 to 3, and the sum of the numbers is 28. What is the smaller number? 5 6 7 8

OpenStudy (anonymous):

Multiply: 7 * 1 7 * 3

OpenStudy (anonymous):

7*1= 7 7*3= 21

OpenStudy (anonymous):

Wait, that is all you have to do?

OpenStudy (anonymous):

Can you do that? @Dlwillis

OpenStudy (anonymous):

Yeah. Lol, and 21 + 7 is 28. I feel so stupid now. xD thank you!

OpenStudy (anonymous):

Oh, nvm..yes so we have 7/21 7 + 21 = 28.

OpenStudy (anonymous):

\[\frac{ x }{ y }=\frac{ 1 }{ 3 },y=3 x\] x+y=28 x+3x=28 4x=28 x=7 y=?

OpenStudy (anonymous):

So your answer should be 7..

OpenStudy (anonymous):

Sorry, OS was lagging for a minute there.

OpenStudy (anonymous):

Oh, yeah. Lol, i hate when it lags!

OpenStudy (bradely):

a/b = 1/3 b=3a a+b =28 a+3a =28 4a = 28 a =7 Small number is 7 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

y=28-7=21

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

can one of you guys please help me please i am dsads i just made up that name please help me

OpenStudy (anonymous):

What do you need help with?